OpenStudy (kikuo):

I am learning about scientific notation, and my book said... -when multiplying numbers in scientific notation multiply the numbers before the multiplication sign and add the exponent -for example (3.0x10^8m/s)x(5.0x10^2s)=15x10^10=1.5x10^11m However, I was confused why it was 15x10^10 and not 15x100^10 since 10x10=100.

OpenStudy (jtug6):

It's because you have a common base of 10, so when you multiply them together youre essentially adding the exponents, just as you would to any other common base, such as x^5 * x^3. It'd be x^8 and you wouldn't account for x*x (or 10*10 in this case). So: (3 * 10^8m/s) * (5 * 10^2s) = 15*10^10m or you can move divide 10 by 15 and then add 1 to the exponent to account for the division, so 1.5*10^11m. Hope that clears things up ask me if you're still stuck!

OpenStudy (kikuo):

Makes sense! But what's a common base? @jtug6

OpenStudy (jtug6):

It's the bottom number that is raising some power. So like, here are some bases: 5^10, 5 is the base x^19, x is the base 2.54x10^19, 10 is the base

OpenStudy (kikuo):

Why is common base used here but not when we combine like terms? For eg when I was in my algebra 1 class (if I remember correctly) x^3*x^3=something around 2x^3 (not sure if three is supposed to be three or 9) Though, point is, what makes this different from combining them? How do I know when to apply common base?

OpenStudy (jtug6):

Uhhh well if the answer was 2x^3 or 2x^9 it'd still be incorrect. x^3 * x^3 is x^6 because you have that x term in common. So when you're multiplying two common bases, you add their exponents as a result. So, x^3 * x^3 is x^(3 + 3) = x^6

ganeshie8 (ganeshie8):

x^3 is a short form for x*x*x

OpenStudy (jtug6):

ganeshie might have a better job of explaining it, too. ^^ lol

ganeshie8 (ganeshie8):

x^3 * x^3 = x*x*x * x*x*x How many x's do you see on the right hand side ?

OpenStudy (kikuo):

6 And ah so this only applies if you have exponents?

ganeshie8 (ganeshie8):

Yes, try to see the big picture, its really easy.

ganeshie8 (ganeshie8):

For example, if you want to multiply 10 by itself 8 times. We express it as : 10 * 10 * 10 * 10 * 10 * 10 * 10 * 10

ganeshie8 (ganeshie8):

That's too painful and waste of time and space

ganeshie8 (ganeshie8):

That's one reason we invented the idea of an exponent. That crazy multiplication is expressed in compact form as : 10^8

ganeshie8 (ganeshie8):

Do you appreciate how neat `10^8` looks compared to `10 * 10 * 10 * 10 * 10 * 10 * 10 * 10` ?

ILovePuppiesLol (ilovepuppieslol):

wow u guys are really good at this math thing, Im trying to do geometry and its really hard

ganeshie8 (ganeshie8):

Geometry is indeed hard. But I think you will do great if you know how to use compass and straightedge efficiently...

ILovePuppiesLol (ilovepuppieslol):

yes i just started today and i was doing like segment bisectors with them

ganeshie8 (ganeshie8):

Heard of euclidea game ?

OpenStudy (kikuo):

Yes. It's much easier. Why do you only keep the base the same after multiplication if two bases are the same?

ganeshie8 (ganeshie8):

Let's work the given example and see

ganeshie8 (ganeshie8):

You want to simplify (3.0x10^8m/s)x(5.0x10^2s)

ganeshie8 (ganeshie8):

(3.0x10^8)x(5.0x10^2) is same as (3.0 x 5.0) x (10^8 x 10^2)

ILovePuppiesLol (ilovepuppieslol):

yeah kai plays it, he showed me it

ganeshie8 (ganeshie8):

With me so far ?

ganeshie8 (ganeshie8):

Its a nice way to learn geometry http://www.euclidea.xyz/en/game/#/packs/Alpha

ILovePuppiesLol (ilovepuppieslol):

yeah, ill check it out, but the tools are too hard to use

OpenStudy (kikuo):

Yes

ganeshie8 (ganeshie8):

(3.0x10^8)x(5.0x10^2) is same as (3.0 x 5.0) x (10^8 x 10^2) (15) x (10^8 x 10^2)

ganeshie8 (ganeshie8):

Let's look at (10^8 x 10^2) closely

ganeshie8 (ganeshie8):

10^8 is a short form for 10*10*10*10* 10*10*10*10 10^2 is a short form for 10*10 therefore, 10^8 x 10^2 = 10*10*10*10* 10*10*10*10 x 10*10 How many 10's do you see on the right hand side ?

OpenStudy (kikuo):

10

ganeshie8 (ganeshie8):

Yeah you will need to get used to the online tools, then only you will enjoy the game :)

ganeshie8 (ganeshie8):

10^8 x 10^2 = 10*10*10*10* 10*10*10*10 x 10*10 = 10^10

ganeshie8 (ganeshie8):

Notice also that 8+2 = 10. Can we say 10^8 x 10^2 = 10^(8+2) ?

OpenStudy (kikuo):

Yes. Though, I'm curious, why does not adding common bases make sense? And how does 10^10 turn into 10^11?

ganeshie8 (ganeshie8):

We don't need to know what a base is here. We are doing series of multiplications here, thats all.

ganeshie8 (ganeshie8):

(3.0x10^8)x(5.0x10^2) is same as (3.0 x 5.0) x (10^8 x 10^2) (15) x (10^8 x 10^2) (15) x (10^10)

ganeshie8 (ganeshie8):

1.5 x 10 = ?

OpenStudy (kikuo):

I mean why does it make sense for 10^8x10^2 to = 10^10 and not 100^10 15

ganeshie8 (ganeshie8):

Let's do it again from scratch

ganeshie8 (ganeshie8):

10^8 = ?

OpenStudy (kikuo):

10x10x10x10x10x10x10x10

ganeshie8 (ganeshie8):

why, why is it not 100x100x100x100x 100x100x100x100 ?

OpenStudy (kikuo):

Because exponents merely tell you how many times the base is going to be multiplied by itself each time

ganeshie8 (ganeshie8):

Good. Keep that in mind for the next few minutes.

ganeshie8 (ganeshie8):

10^2 = ?

OpenStudy (kikuo):

10x10

ganeshie8 (ganeshie8):

10^8 = 10x10x10x10x10x10x10x10 10^2 = 10x10 --------------------------------------------- 10^8 x 10^2 = ?

ganeshie8 (ganeshie8):

multiply the right hand sides

OpenStudy (kikuo):

OH I see why it doesn't make sense. Okay, I understand why we don't multiply the bases. 10^10 So where does 10^11 come from?

ganeshie8 (ganeshie8):

All you have to remember is the definition : a^n = a*a*a*... n times

ganeshie8 (ganeshie8):

The rules like a^m * a^n = a^(m+n) and everything else follow from that definition

ganeshie8 (ganeshie8):

If you remember just the rules and ignore the definition, you will be confused and get stuck everywhere.

OpenStudy (kikuo):

(3.0x10^8m/s)x(5.0x10^2s)=15x10^10=1.5x10^11m Alright that makes sense, but in this problem, 15 turns into 1.5 and 10^10 turns into 10^11? Why?

ILovePuppiesLol (ilovepuppieslol):

good night guys

ganeshie8 (ganeshie8):

Look at 15x10^10

ganeshie8 (ganeshie8):

have good sleep @ILovePuppiesLol

ganeshie8 (ganeshie8):

1.5 x 10^1 = ?

OpenStudy (kikuo):

15, but why did it change into 1.5? @ganeshie8

OpenStudy (zzr0ck3r):

15=1.5*10 150=1.5*10^2 1500=1.5*10^3

OpenStudy (kikuo):

Ah, I get it that part now. @zzr0ck3r Why did it turn from 1.5*10^11 instead of 1.5*10^10?

ganeshie8 (ganeshie8):

What has turned to 1.5*10^11 ?

OpenStudy (kikuo):

(3.0x10^8m/s)x(5.0x10^2s)=15x10^10=1.5x10^11m See how it ends with 10^11 Why? This is how it was written in my book btw.

ganeshie8 (ganeshie8):

15x10^10 first replace 15 by 1.5 x 10^1

ganeshie8 (ganeshie8):

what do you get ?

OpenStudy (kikuo):

1.5x10^10

OpenStudy (kikuo):

You there? @ganeshie8

jhonyy9 (jhonyy9):

sorry kikuo but i not can answering you just in this style ,here in what you need help ?

jhonyy9 (jhonyy9):

do you need help in this posted question again ? i think you will can understanding this multiplication of scientific notations with and other example - how you think ?

jhonyy9 (jhonyy9):

an other example

jhonyy9 (jhonyy9):

3*10^2 * 5*10^5 =?

jhonyy9 (jhonyy9):

or an other one (2*10^2) * (3*10^8) = ? for your more easy understanding i separated with parentheses

OpenStudy (kikuo):

I want to understand why it changed from 10^10 to 10^11 Why did that happen? I want to understand specifically that. : )

jhonyy9 (jhonyy9):

@ganeshie8 this 15 *10^10 we transform it in 1,5 *10^11 bc. the scientific notation is in this form just one integer and how many much more the 10 on the necessary power so than you see 15 is formed from two integer numbers and in this way we can transform it in 1,5 and hence the 10 power will be changed in 11 hope this is understandably for you and helped you

jhonyy9 (jhonyy9):

Kikuo do you understand it now ?

OpenStudy (kikuo):

Not quite

jhonyy9 (jhonyy9):

ask me what no

jhonyy9 (jhonyy9):

do you need an other example ?

jhonyy9 (jhonyy9):

2*10^3 *5*10^2 = ?

OpenStudy (kikuo):

I'm not understanding why the 10 turned into 11. Like, I don't understand.

OpenStudy (phi):

\[ 15 \cdot 10^{10}=1.5 \cdot 10^{11} \] why? the rule is that we want the number "out front" to be less than 10 15 is bigger than 10, so we have to tweak it. up above, people showed you \[ 15 = 1.5 \cdot 10^1 \] do you see how that is true ? anyway, if you replace 15 in \[ 15 \cdot 10^{10}\] with this new way of writing 15 as 1.5 * 10^1 you get\[ 1.5 \cdot 10^1 \cdot 10^{10} \] now figure out what \[10^1 \cdot 10^{10} \] is (you should get \( 10^{11}\)) and the answer is \[ 1.5 \cdot 10^{11} \] if this is not clear, please post another question, and people will help some more.